OpenStudy (anonymous):

The direct distance from a starting point to a finish line is 15 miles. Unfortunately, you can't take the direct route. If you travel 12 miles east, how many miles north must you travel to reach the finish line?

OpenStudy (anonymous):

 pythagoras right triangle formula a^2 + b^2 = c^2 x^2 + 12^2 = 15^2